OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (hoblos):

slope-intercept form has a general form of y=ax+b so you need to change your equation to this form y − 4 =3/4 (x + 8) distribute the (3/4) y-4 = (3/4)x + 6 add 4 to both sides y= (3/4)x +10